OpenStudy (anonymous):

Please help with 1 equation

HanAkoSolo (jamierox4ev3r):

same thing here.. just distribute Ex: -2x+4=8x-6 can you figure out the rest? @trollface

OpenStudy (anonymous):

-2x + 4 = 2(4x - 3) - 3(-8 + 4x) -2x + 4 = 8x - 6 +24 -12x -2x + 4 = 8x -12x+24- 6 -2x + 4 = -4x+18 4x-2x=18-4 2x= 14 x= 14/2 x=7 Hence option A is the correct answer. @trollface

OpenStudy (anonymous):

I think so, thanks both of yous!

HanAkoSolo (jamierox4ev3r):

no problem :) and for the future @dpasingh , try not to solve the problem for the user. let them figure it out, but thanks for helping, keep up the great work